• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 11:16:41 on September 27, 2010

    Entry number 332914

    keyword=Experiment status update/summary

    Mon Sep 27 11:14:56 EDT 2010
    - See this link for a runplan outline
    - MCC/Accelerator
      - 'Hot checkout' of beamline magnets ONLY is deferred until Tuesday at
        - PD will check with Ron Lauze as to whether this early partial
    checkout is
          worth the overhead
      x delivery to Hall B should happen today
        - note problem with CHL will shut things down tomorrow
      x some issues with cryo levels in RF zones over the weekend are to be
        investigated (CHL contamination problem?)
      x Harps, PSS tested ready to go
        x no "show stoppers" on MCC side other than cryo-related problems
    - ESR
      x warmed up over weekend, cleared (they hope) the ice, recooling today
    - CHL has contamination
      - 1--2 day down minimum, starting Tuesday morning
      - not sure how/if this impacts ESR cooldown
      - meeting on CHL issue at 1:30pm today
    - Other Cryo-related issues:
      * Note that cooling the LHRS magnets will need to be done with Hall open
        - will take about a week to cool the magnets, 
        - my understanding is that we should be able to take some beam in the
          evenings in parallel with this process, but we will have to work
    with the
          Techs on scheduling
      * Note that cooling target will need to be done with Hall open
      * Note that cooling Moller will need to be done with Hall open (just
        on this one)
    - Target (Meekins)
      x spoke with Folts and Meekins
      - leak not found yet, still investigating this morning
      - in afternoon, Meekins will have to work on Qweak target...
    - Arc Energy system (Arun)
      - 9th dipole is functioning
      - scanner HW and SW status unknown - emailed Arun requesting status update
    - Lumis (Kalyan)
      x mostly ready, but noisy
        x is connected to RHRS scalers and injected using event 140 system
        - data aren't exposed to the software yet (not sure what the problem is)
      x lots of noise on system (ringing and 60Hz)
        - adding RC filter !?
    - Moller (Sasha, Jixie)
      x ready to go, tested at 1kHz MPS --> HW and SW working
      x tested system with LED pulser, in good shape
      x magnet is already precooled using LN2, just waiting on ESR now
        - should be able to cool in a day or so
        - Note that cooling magnet will need to be done with Hall open (I
      - Note that we will eventually need to do alignment through solenoid, must
        bypass compton chicane for this
        - could impact scheduling of beam quality tuning
          - ie. do Moller alignment first, then steer and optimize through
    - Compton (Eric Fuchey)
      x DAQ mostly ready
        - missing 'prescaler' <-- Alex
        - Megan Fried is the only person who really understands the DAQ software
          - she is willing/able to come onsite for 1--2 weeks, but this needs to
            be scheduled
      x HW 'should' be OK, but is is not powered on yet
      - Eric will work with others to test system and get back to me
    - DAQ (Alex, others)
      - 'Logic module' needs to be updated/fixed so that LHRS and Calorimeter
        singles can be taken
      - there is discussion about requiring all trigs but T5 to be disabled
    for the
        production configuration since the Calorim ADCs don't get gated on an
        - wasn't clear to me if the 'logic module' fix above solves this, or if
          there really does need to be multiple CODA configurations to handle
          singles triggers...
      - was problem with 1875 TDC in LHRS as of Friday
        - fixed, replace now?
    - BPMs, BCMs, Raster status
      - all 'should' be plugged in somewhere, but not fully verified on
    software side
        - BCM scalers are working
        - need to verify BPMs plugged into ADCs (in both arms!)
        - need to verify Raster signals plugged into ADCS (in both arms)
    - Calorimeter (Alex)
      x ARS unit(s) had firmware updated over weekend
        - seems to have fixed a crash problem
      - some readout/sync errors remain
      - needs to be incorporated into complete DAQ (LHRS + Calorimeter) so it is
        possible to take singles
        - at the moment the Calorim ADCs only get a gate on a Calorim
    trigger, any
          other trigger will cause a readout problem
    - LHRS (Isham, Julie)
      - all detectors were working at one point, but haven't been tested for a
        - can't take LHRS singles cosmics triggers in present configuration
      - EDTM status unknown
        - not clear if experiment cares about it
      - [Segal] NOTE that the power supplies have seem major repairs and need
    to be
        recommissioned.  This will take 2--3 day shifts _after_ the magnets
    are cold.